# On Geometry-Enhanced Parameter-Efficient Fine-Tuning for 3D Scene Segmentation

### v1

@@ 3.2 Our Methodology: Geometry Encoding Mixer (GEM)
where $\bm{x}$ denotes input points and $\text{pos}(\cdot)$ is the pre-trained positional embedding. Fig. 2 illustrats the overall structure.
@@
With a common kernel dimension $k=3$, the spatial adapter touches at most $k^{3}$ neighbours per point and adds $2rd+k^{3}r^{2}$ parameters, rendering a complexity of $\mathcal{O}(2ndr+nk^{3}r^{2})=\mathcal{O}(nd)$ given $k,r\ll n,d$. It thus functions as an efficient convolution-based positional encoding [11, 29] in parallel to the pre-trained positional encoding, thereby capturing fine spatial layout to match the target distribution.
@@
As both attention costs $O(nm)$ with $m\ll n$, global aggregation is affordable and at the same level of complexity as the conventional prompt tuning [37, 36]. Furthermore, we update $\bm{L}\leftarrow\bm{L}+\bm{L}_{c}$ at each adapter insertion to share across layers, yielding dynamic prompts that capture the context of current scene, unlike static prefixes in prompt tuning.
@@ Discussion.
In comparison to existing PEFT methods, GEM explicitly biases the adaptation to 3D geometry and context. The spatial adapter captures localized variations that generic adapters, such as LoRA and basic adapters, would miss, since those ignore spatial structure and use the positional encodings from pre-training. The context adapter circumvent the local attention design to provide scene-wide context for efficient adaptation with few parameters. By combining these two components, our approach enables scene-aware adaptation of 3D transformers, where each point is tuned with respect to both its neighbors and the entire point cloud. It thus results in significantly improved performance on 3D scene datasets with only a lightweight adaptation overhead.
@@ 4 Experiments
We present the results of our proposed GEM on semantic segmentation of large-scale 3D scene datasets and experiment with both self-supervised pre-trained backbone, Sonata [74], and another one, PTv3-PPT [77], for investigation. We also provide ablation studies to better reveal the detailed effects of different components.
@@ 4.1 Experimental Setup
We primarily follow the comprehensive protocols proposed in Sonata [74], covering ScanNet [12], ScanNet200 [60], ScanNet++ [82] and S3DIS [1].
We adopt two representative pre-trained backbones, the Sonata model [74] with large-scale self-supervised training and the PTv3-PPT [77] with supervised pre-training on multiple large-scale curated datasets. We compare GEM with the existing popular PEFT methods, including bias-tuning from BitFit [5], Adapter [25], LoRA [26], and Prompt Tuning [37]. We consider linear probing as a baseline and the full fine-tuned model as our target strong reference. In addition, we note that the Sonata (full.) specifically introduce a task-specific decoder for semantic segmentation during the fine-tuning, which may not reflect the fair comparison. Therefore, we revive to the standard full fine-tuning without additional decoder network, which is denoted by Sonata (ft.).
For training, we follow the widely accepted fine-tuning setups to update only the inserted or selected weights with the pre-trained backbone weights remaining frozen. All PEFT baselines follow the implementations from released code, adopt the suggested common practice [47, 52], and are tuned to their best validation setting in Fig. 1(c). Specifically, We set the default rank to be $r=32$ and the number of learnable tokens to be $m=4$. More details are given in the supplementary.
@@ 4.2 Performance Comparison
Semantic Seg. Params ScanNet Val [12] ScanNet200 Val [60] ScanNet++ Val [82] S3DIS Area 5 [1] S3DIS 6-fold [1]
@@
+ BitFit [5] 0.2M 0.2% 74.7 84.7 90.8 32.5 45.5 82.0 42.4 56.5 85.7 73.9 82.0 91.5 76.1 87.1 91.0
+ Adapter [25] 2.8M 2.5% 77.0 85.4 91.8 33.6 45.7 82.5 42.6 57.5 85.6 73.8 82.9 91.5 76.4 87.5 91.4
+ LoRA [26] 1.9M 1.7% 76.7 85.7 91.7 33.6 45.5 82.7 44.2 58.0 86.5 74.5 83.2 91.5 77.4 87.8 91.2
+ Prompt Tunning [37] 5.5M 4.8% 74.3 84.1 90.5 31.4 44.4 81.6 41.2 56.2 84.8 73.4 82.5 91.0 73.7 86.5 90.5
@@
Tab. 1 shows that GEM consistently surpasses all representative PEFT methods across datasets. It matches the performance of full fine-tuning on most datasets and even exceeds it on ScanNet++[82]. ScanNet++ comprises large and diverse scenes captured at sub-millimeter resolution, diverging greatly from the spatial distribution of common pre-training datasets such as ScanNet [12]. Under this domain gap, the joint modeling of local spatial cues and global scene context introduced by GEM proves especially beneficial.
Interestingly, LoRA [26] and adapter [25] deliver similar scores, despite acting on different transformer components. This observation implies that, when updates remains local and geometric is not modeled explicitly, the precise choice of adaptation target can be secondary. In addition, prompt tuning [37] underperforms even the linear probing baseline in the S3DIS 6-fold evaluation, revealing the penalty of ignoring spatial structure during fine-tuning.
@@
We assess PEFT methods under the scenarios of limited data and annotations in Tab. 3. The results demonstrate the exceptional data efficiency of GEM, which outperforms both other PEFT methods and the full fine-tuning counterparts. Notably, under extreme data scarcity, such as 1% of the labeled scenes and limited annotations (20 points per scene), GEM surpasses both Sonata (full.) and Sonata (ft.), highlighting its superiority in low-data regimes.
@@
While self-supervised methods dominate in the language domain, supervised pre-training remains competitive for vision [33]. To probe the limits of PEFT under large-scale 3D supervision, we adopt recent advances in 3D supervised pre-training [77], which achieve leading performance by training on a large collection of curated, labeled scene datasets. As shown in Tab. 3, existing PEFT methods can even degrade performance, likely suffering from negative transfer [86, 72]. In contrast, GEM improves upon supervised backbone, further outperforming the larger PTv3 [75] with only 1.6% additional parameters.
@@
A dedicated segmentation decoder ( 13% of the total parameters) is known to lift fully fine-tuned baselines. We therefore ask whether PEFT can still add value in the presence of such a task-specific head.
Table 5 answers in the affirmative: GEM outperforms all competing PEFT variants and even surpasses its own fully fine-tuned counterpart, setting a new state-of-the-art on ScanNet. The result underscores the merit of jointly modeling local geometry and scene-level context, even when the decoder already provides task-specific capacity.
@@
During the original evaluation of Sonata [74], it employs separate pre-trained models for indoor and outdoor scenarios, due to the significant domain gaps between indoor and outdoor scenes. Here we test a stronger setting: transferring an indoor pre-trained backbone to an outdoor autonomous-driving benchmark. We replace the input layer with a randomly initialized counterpart to match dimensionality and freeze all remaining weights.
Table 5 shows that GEM narrows the gap to a model trained from scratch on outdoor data, yet a noticeable performance gap persists. Closing this gap remains an interesting direction for future work.
[…]
@@
+ BitFit [5] 46.5 64.9 69.9 71.7 74.7 71.0 72.3 72.9 73.6 74.7
+ Adapter [25] 46.4 64.2 70.1 72.2 77.0 71.8 73.4 74.7 75.0 77.0
+ LoRA [26] 46.6 63.0 70.1 72.6 76.7 72.1 73.6 75.2 75.5 76.7
+ Prompt Tunning [37] 45.5 62.6 68.9 71.1 74.3 70.2 71.5 72.5 72.8 74.3
@@
Supervised Pre-train. Params ScanNet Val [12]
@@
+ BitFit [5] 0.2M 0.2% 78.2 85.6 92.3
+ Adapter [25] 1.8M 1.8% 78.5 85.9 92.4
+ LoRA [26] 1.4M 1.7% 78.4 86.0 92.4
+ Prompt Tunning [37] 4.8M 4.8% 78.3 85.9 92.4
@@
[…]
with Decoder Params ScanNet Val [12]
@@
+ BitFit [5] 16.4M 13.2% 78.1 86.1 92.3
+ Adapter [25] 19.1M 15.0% 78.2 86.5 92.3
+ LoRA [26] 18.2M 14.3% 78.9 87.0 92.5
+ Prompt Tunning [37] 22.6M 17.2% 77.4 86.1 92.1
@@
Outdoor Seg. Params Sem.KITTI Val [4]
@@
+ BitFit [5] 0.2M 0.2% 59.9 70.6 91.0
+ Adapter [25] 2.8M 2.5% 62.5 72.2 92.0
+ LoRA [26] 1.9M 1.7% 63.8 74.0 92.0
+ Prompt Tunning [37] 5.5M 4.8% 59.1 71.0 90.8
@@ 4.3 Ablations and Analsys
We mainly consider the linear probing, Sonata (lin.) as baseline and ablate on ScanNet [12] to better investigate the key factprs for effective PEFT in 3D scenes, shown in Tab. 6. For more studies on the effect of available parameters, please refer to the supplement.
@@
To validate our initial hypothesis, we study the individual effectiveness of the SA and CA in LABEL:tbl:abl_comp. We find that both adapters can can already be superior to the baseline and competitive to the existing PEFT methods. We notice that the SA takes the majority part of the introduced paramteres, which is partially due to the inherently expensive 3D convolutoin with $k^{3}$ kenrels. Such challenge is also 3D specific and also urges us to develop the CA, rather than soly relying on the positional encoding to capture 3D geometry. By combining the two form of spatial adapters, we obtain GEM that reaches the best performance by comprehensively exploring the 3D scene while remaining parameter-efficient.
@@
In LABEL:tbl:abl_tokens, we are motivated to further investigate how efficient and effective of the proposed CA could be. We thus study how few tokens it could afford before failing in capturing the global context and obtaining a similar performance to SA-only. Surprisingly, we find that CA can use as few as one single token to effectively express the global scene context, and obtaining clear improvement over SA-only approach. We consider this to be the effectiveness of weight decomposition, where we are akin to approximating the $n^{2}$ global self-attention weights with two $n\times 1$ matrix. In comparison to LoRA that decompose the in weight space, we decompose the spatial attention weight matrices and form a spatial bottleneck at token dimension. In addition, while we are aware of such similar topic has been similarly explored for efficient attentions [30, 35], we are, to the best of our knowledge, the first to motivate their effective application as PEFT methods for 3D scenes understanding.
@@
The last experiment, LABEL:tbl:abl_dist probes whether the latent tokens learned by CA should remain layer-local or be shared across the hierarchy, assessing its ability of capturing dynamic and scene-specific context. When each layer keeps its own bank of tokens (N/A), GEM already surpasses SA-only, confirming that CA supplies complementary global cues. Re-using the same tokens within each encoder stage (per-stage) yields a further yet modest gain, suggesting that a coherent context inside each resolution level helps the network reconcile local geometry with mid-range structure.
The strongest result emerges when a single set of latent tokens are shared by all transformer blocks. We attribute this improvement to its coincidence with the fact that 3D point cloud represent the same underlying 3D geometry in sipte of its irregularity and sparsity. From this perspective, such globally shared latent tokens confines the model to adapt to the true underlying scene, rather than overfitting on irrelavent pattern due to noisy sampled point cloud.
@@
With the above assumption on latent tokens regularize the model to capture the underlying 3D scene, we visualize its attentional weights and compare to other similar PEFT method that can also inject global context, such as the prompt tuning.
As shown in Fig. 3, starting from the first stage, our latent tokens learn to produce much crisper attention weights following the geometric cues, such as the simple geometric primitives of floor surface. In deeper stages, latent tokens start to pick up more salient geometric objects, such as clutter on the table and sofa. In contrast, the static tokens introduced by prompt tuning produce attention blurry weights that span accross the boundaries of objects and scene areas. As a result, it is suggested that the proposed latent tokens can indeed deliver scene-wide context that broadcast the geometry cues. Together with the local spatial structure extracted by 3D convolutions, it thus results in its superior performance as an effective PEFT methods for 3D scenes.
@@
Table 6: Ablations on GEM. If not specified, the backbone is the self-supervised pre-trained Sonata [74] with no additional task-specific decoder, evaluated on ScanNet [12]. Default settings are marked in gray.

### v2

@@ 3.2 Our Methodology: Geometry Encoding Mixer (GEM)
where $\bm{x}$ denotes input points and $\text{pos}(\cdot)$ is the pre-trained positional embedding. Fig.˜2 illustrats the overall structure.
@@
With a common kernel dimension $k=3$, the spatial adapter touches at most $k^{3}$ neighbours per point and adds $2rd+k^{3}r^{2}$ parameters, rendering a complexity of $\mathcal{O}(2ndr+nk^{3}r^{2})=\mathcal{O}(nd)$ given $k,r\ll n,d$. It thus functions as an efficient convolution-based positional encoding [pos_cpe, pos_conv] in parallel to the pre-trained positional encoding, thereby capturing fine spatial layout to match the target distribution.
@@
As both attention costs $O(nm)$ with $m\ll n$, global aggregation is affordable and at the same level of complexity as the conventional prompt tuning [prefix, prefix_prompt]. Furthermore, we update $\bm{L}\leftarrow\bm{L}+\bm{L}_{c}$ at each adapter insertion to share across layers, yielding dynamic prompts that capture the context of current scene, unlike static prefixes in prompt tuning.
@@ Discussion.
In comparison to existing PEFT methods, GEM explicitly biases the adaptation to 3D geometry and context. The spatial adapter captures localized variations that generic adapters, such as LoRA and basic adapters, would miss, since those ignore spatial structure and use the positional encodings from pre-training. The context adapter circumvents the local attention design to provide scene-wide context for efficient adaptation with few parameters. By combining these two components, our approach enables scene-aware adaptation of 3D transformers, where each point is tuned with respect to both its neighbors and the entire point cloud. It thus results in significantly improved performance on 3D scene datasets with only a lightweight adaptation overhead [^1] We study the empirical overheads in the supplementary Appendix A. .
@@ 4 Experiments
We present the results of our proposed GEM on semantic segmentation of large-scale 3D scene datasets and experiment with both self-supervised pre-trained backbone, Sonata [sonata], and supervised pretrained one, PTv3-PPT [ppt], for investigation [^2] For more generalization and comparisons in other settings, such as convolutional model and 3D shape analysis, please refer to the supplementary Appendix B. . We also provide ablation studies to reveal the detailed effects of different components better.
@@ 4.1 Experimental Setup
We primarily follow the comprehensive protocols proposed in Sonata [sonata], covering ScanNet [scannet], ScanNet200 [scannet200], ScanNet++ [scannetpp] and S3DIS [s3dis].
We adopt two representative pre-trained backbones, the Sonata model [sonata] with large-scale self-supervised training and the PTv3-PPT [ppt] with supervised pre-training on multiple large-scale curated datasets. We compare GEM with the existing popular PEFT methods, including bias-tuning from BitFit [select_bitfit], Adapter [adap], LoRA [lora], and Prompt Tuning [prefix]. We consider linear probing as a baseline and the full fine-tuned model as our target strong reference. In addition, we note that the Sonata (full.) specifically introduces a task-specific decoder for semantic segmentation during the fine-tuning, which may not reflect the fair comparison. Therefore, we revive to the standard full fine-tuning without additional decoder network, which is denoted by Sonata (ft.).
For training, we follow the widely accepted fine-tuning setups to update only the inserted or selected weights with the pre-trained backbone weights remaining frozen. All PEFT baselines follow the implementations from released code, adopt the suggested common practice [code_peft, code_adapters], and are tuned to their best validation setting in Fig.˜1(c). Specifically, we set the default rank to be $r=32$ and the number of learnable tokens to be $m=4$. More details are given in the supplementary.
@@ 4.2 Performance Comparison
Semantic Seg. Params ScanNet Val [scannet] ScanNet200 Val [scannet200] ScanNet++ Val [scannetpp] S3DIS Area 5 [s3dis] S3DIS 6-fold [s3dis]
@@
+ BitFit [select_bitfit] 0.2M 0.2% 74.7 84.7 90.8 32.5 45.5 82.0 42.4 56.5 85.7 73.9 82.0 91.5 76.1 87.1 91.0
+ Adapter [adap] 2.8M 2.5% 77.0 85.4 91.8 33.6 45.7 82.5 42.6 57.5 85.6 73.8 82.9 91.5 76.4 87.5 91.4
+ LoRA [lora] 1.9M 1.7% 76.7 85.7 91.7 33.6 45.5 82.7 44.2 58.0 86.5 74.5 83.2 91.5 77.4 87.8 91.2
+ Prompt Tunning [prefix] 5.5M 4.8% 74.3 84.1 90.5 31.4 44.4 81.6 41.2 56.2 84.8 73.4 82.5 91.0 73.7 86.5 90.5
@@
Tab.˜1 shows that GEM consistently surpasses all representative PEFT methods across datasets. It matches the performance of full fine-tuning on most datasets and even exceeds it on ScanNet++[scannetpp]. ScanNet++ comprises large and diverse scenes captured at sub-millimeter resolution, diverging greatly from the spatial distribution of common pre-training datasets such as ScanNet [scannet]. Under this domain gap, the joint modeling of local spatial cues and global scene context introduced by GEM proves especially beneficial.
Interestingly, LoRA [lora] and adapter [adap] deliver similar scores, despite acting on different transformer components. This observation implies that, when updates remain local and geometric is not modeled explicitly, the precise choice of adaptation target can be secondary. In addition, prompt tuning [prefix] underperforms even the linear probing baseline in the S3DIS 6-fold evaluation, revealing the penalty of ignoring spatial structure during fine-tuning.
@@
We assess PEFT methods under the scenarios of limited data and annotations in Tab.˜3. The results demonstrate the exceptional data efficiency of GEM, which outperforms both other PEFT methods and the full fine-tuning counterparts. Notably, under extreme data scarcity, such as 1% of the labeled scenes and limited annotations (20 points per scene), GEM surpasses both Sonata (full.) and Sonata (ft.), highlighting its superiority in low-data regimes.
@@
While self-supervised methods dominate the language domain, supervised pre-training remains competitive for vision [sam]. To probe the limits of PEFT under large-scale 3D supervision, we adopt recent advances in 3D supervised pre-training [ppt], which achieve leading performance by training on a large collection of curated, labeled scene datasets. As shown in Tab.˜3, existing PEFT methods can even degrade performance, likely suffering from negative transfer [nt_deep, nt_avoid]. In contrast, GEM improves upon supervised backbone, further outperforming the larger PTv3 [ptv3] with only 1.6% additional parameters.
@@
A dedicated segmentation decoder (13% of the total parameters) is known to lift fully fine-tuned baselines. We therefore ask whether PEFT can still add value in the presence of such a task-specific head. Tab.˜5 answers in the affirmative: GEM outperforms all competing PEFT variants and even surpasses its own fully fine-tuned counterpart, setting a new state-of-the-art on ScanNet. The result underscores the merit of jointly modeling local geometry and scene-level context, even when the decoder already provides task-specific capacity.
@@
During the original evaluation of Sonata [sonata], it employs separate pre-trained models for indoor and outdoor scenarios, due to the significant domain gaps between indoor and outdoor scenes. Here, we test a stronger setting: transferring an indoor pre-trained backbone to an outdoor autonomous-driving benchmark. We replace the input layer with a randomly initialized counterpart to match dimensionality and freeze all remaining weights.
Tab.˜5 shows that GEM narrows the gap to a model trained from scratch on outdoor data, yet a noticeable performance gap persists. Closing this gap remains an interesting direction for future work.
@@
+ BitFit [select_bitfit] 46.5 64.9 69.9 71.7 74.7 71.0 72.3 72.9 73.6 74.7
+ Adapter [adap] 46.4 64.2 70.1 72.2 77.0 71.8 73.4 74.7 75.0 77.0
+ LoRA [lora] 46.6 63.0 70.1 72.6 76.7 72.1 73.6 75.2 75.5 76.7
+ Prompt Tunning [prefix] 45.5 62.6 68.9 71.1 74.3 70.2 71.5 72.5 72.8 74.3
@@
Supervised Pre-train. Params ScanNet Val [scannet]
@@
+ BitFit [select_bitfit] 0.2M 0.2% 78.2 85.6 92.3
+ Adapter [adap] 1.8M 1.8% 78.5 85.9 92.4
+ LoRA [lora] 1.4M 1.7% 78.4 86.0 92.4
+ Prompt Tunning [prefix] 4.8M 4.8% 78.3 85.9 92.4
@@
with Decoder Params ScanNet Val [scannet]
@@
+ BitFit [select_bitfit] 16.4M 13.2% 78.1 86.1 92.3
+ Adapter [adap] 19.1M 15.0% 78.2 86.5 92.3
+ LoRA [lora] 18.2M 14.3% 78.9 87.0 92.5
+ Prompt Tunning [prefix] 22.6M 17.2% 77.4 86.1 92.1
@@
Outdoor Seg. Params Sem.KITTI Val [semkitti]
@@
+ BitFit [select_bitfit] 0.2M 0.2% 59.9 70.6 91.0
+ Adapter [adap] 2.8M 2.5% 62.5 72.2 92.0
+ LoRA [lora] 1.9M 1.7% 63.8 74.0 92.0
+ Prompt Tunning [prefix] 5.5M 4.8% 59.1 71.0 90.8
@@ 4.3 Ablations and Analsys
We mainly consider the linear probing, Sonata (lin.), as baseline and ablate on ScanNet [scannet] to better investigate the key factors for effective PEFT in 3D scenes, shown in Tab.˜6. For more studies on the effect of available parameters, please refer to the supplement.
@@
To validate our initial hypothesis, we study the individual effectiveness of the SA and CA in LABEL:tbl:abl_comp. We find that both adapters can already be superior to the baseline and competitive to the existing PEFT methods. We notice that the SA takes the majority part of the introduced parameters, which is partially due to the inherently expensive 3D convolution with $k^{3}$ kernels. Such challenge is also 3D specific and urges us to develop the CA, rather than solely relying on the positional encoding to capture 3D geometry. By combining the two form of spatial adapters, we obtain GEM that reaches the best performance by comprehensively exploring the 3D scene while remaining parameter-efficient.
@@
In LABEL:tbl:abl_tokens, we are motivated to further investigate how efficient and effective the proposed CA could be. We thus study how few tokens it could afford before failing to capture the global context. Surprisingly, we find that CA can use as few as one single token to effectively express the global scene context, obtaining clear improvement over SA-only approach. We consider this to be the effectiveness of weight decomposition, where we are akin to approximating the $n^{2}$ global self-attention weights with two $n\times 1$ matrices. In comparison to LoRA that decomposes in the weight space, we decompose the spatial attention weight matrices and form a spatial bottleneck at token dimension. In addition, while we are aware that related topics have been similarly explored for efficient attentions [tr_perceiver, tr_set], we are, to the best of our knowledge, the first to motivate their effective application as PEFT methods for 3D scenes understanding.
@@
The last experiment, LABEL:tbl:abl_dist probes whether the latent tokens learned by CA should remain layer-local or be shared across the hierarchy, assessing its ability to capture dynamic and scene-specific context. When each layer keeps its own bank of tokens (N/A), GEM already surpasses SA-only, confirming that CA supplies complementary global cues. Re-using the same tokens within each encoder stage (per-stage) yields a further yet modest gain, suggesting that a coherent context inside each resolution level helps the network reconcile local geometry with mid-range structure.
The strongest result emerges when a single set of latent tokens is shared by all transformer blocks. We attribute this improvement to its consistency with the fact that 3D point clouds represent the same underlying 3D geometry in spite of their irregularity and sparsity. From this perspective, such globally shared latent tokens confine the model to adapt to the actual underlying scene, rather than overfitting on irrelevant patterns due to the noisy sampled point cloud.
@@
Under the above assumption that latent tokens regularize the model to capture the underlying 3D scene, we visualize the attention weights and compare them to those produced by other PEFT methods that also inject global context, such as prompt tuning.
As shown in Fig.˜3, starting from the first stage, our latent tokens learn to produce crisper attention weights that follow geometric cues, such as the simple geometric primitives of floor surface. In deeper stages, the tokens progressively attend to more salient geometric objects, including clutter on the table and sofa. In contrast, the static tokens introduced by prompt tuning produce yield blurrier attention patterns that often span across object and region boundaries.
These observations suggest that the proposed latent tokens effectively capture and broadcast scene-wide geometric context. Combined with the local spatial structure extracted by 3D convolutions, this leads to superior performance, establishing our method as an effective PEFT approach for 3D scene understanding.
@@
Table 6: Ablations on GEM. If not specified, the backbone is the self-supervised pre-trained Sonata [sonata] with no additional task-specific decoder, evaluated on ScanNet [scannet]. Default settings are marked in gray.